# Translating Images to Road Network: A Sequence-to-Sequence Perspective

### v2

@@ IV-B Implementation
Pretrain We follow LSS [13] for BEV Encoder. The input images are resized to $128\times 352$, and the target BEV area is from -48 to 48m in x-direction (roll) and -32 to 32m in y-direction (pitch) with resolution $1$m in ego coordinate system. ResNet-50 [57] or VoVNetV2 [58] are used as image backbone. The ResNet-50 [57] model is pretrained using DeepLabV3+ [59] on the Cityscapes dataset[60], and the VoVNetV2 is pretrained on FCOS3D [61]. To be noted that the VoVNetV2 is pretrained on default split of nuScenes [24] dataset, therefore, we will not use the VoVNetV2 in non-overfitting split. Initially, we pretrain the BEV encoder on centerline segmentation, following all the training strategy outlined in LSS [13]. During training of RoadNetTransformer, we load and freeze the parameters of BEV encoder.
RoadNetTransformer Details of three variants are shown below. AR-RNTR: The length of RoadNet Sequence is padded to $6\times 100$. Transformer decoder layers of AR-RNTR is set as 6. Our AR-RNTR is trained on 300 epochs with learning rate $2\times 10^{-4}$, batch size as $2\times 8$. During the training process, we apply random flip, random rotation and random scaling on BEV feature similar to [29, 16]. SAR-RNTR: Transformer decoder layers of SAR-RNTR is set as 6 (Key-point) plus 3 (Parallel-seq). We use noise vertices to pad all sub-sequences to the length of $6\times 18$, and the max number of key-points is set to 34. The strategies for data augmentation and training are the same as AR-RNTR. NAR-RNTR: During training, we load parameter of trained SAR-RNTR and then finetune it with mask language modeling strategy for 100 extra epochs. During inference, iteration number is set to 3.
Topology-Inherited Training We apply Topology-Inherited Training only to AR-RNTR. The basic settings of AR-RNTR are the same as those previously introduced for RoadNetTransformer. In the first stage, we train on 20Hz data for 300 epochs. We use more data augmentation than when training RoadNetTransformer. The random rotation is increased from $-45^{\circ}$ to $45^{\circ}$, and the scale is adjusted from 0.9 to 1.1. In the second stage, we train for 24 epochs. In the third stage, we train for 100 epochs.
Topology-Inherited Training We locate Singapore onenorth, Singapore Queenstown, Singapore Hollandvillage, and Boston Seaport from OpenStreetMap [27] and approximately choose the corresponding areas. After exporting the topology data, we retain only the nodes of highway and their secondary type of link roads.
Metrics Thresholds for Mean Precision-Recall and Landmark Precision-Recall are uniformly sampled from $[0.5,5.0]$ m. Thresholds for Reachability Precision-Recall are uniformly sampled from $[0.5,2.5]$ m. We only take into account the reachability between vertices that are within a maximum of 5 edges of connection.
@@ IV-C Results
TABLE I: Comparison of front-camera road network extraction with state of the art on nuScenes [24] PON validation split [25]. ResNet-50 [57] is applied as image backbone by default. M-P, M-R, M-F stand for mean precision/recall/F1-score [17]. Detect stands for Detection ratio metrics [17]. C-P, C-R, C-F stand for connectivity precision/recall/F1-score [17].
@@
Comparison with state of the art We compare our model with previous state-of-the-art methods on high-definition road network topology extraction. To achieve a fair comparison, we only utilize front-view images as input and ResNet-50 [57] trained on ImageNet-1K [63] as backbone. Also, we utilize the PON nuScenes train/val split [25] as also applied in [17, 18]. Table I presents the results of our approach on the nuScenes dataset using the Mean Precision-Recall, Detection, and Connectivity metrics [17]. On the one hand, our model outperforms all three variants across all metrics, demonstrating the superior performance of RNTR in both centerline detection and centerline association estimation. This remarkable improvement on both centerline location and connectivity can be attributed to the unified representation provided by RoadNet Sequence and the global context reasoning capabilities of the Transformer architecture.
On the other hand, unlike in Natural Language Processing, our Semi-Autoregressive and Non-Autoregressive version significantly outperforms the Auto-Regressive version, highlighting our dependency decoupling.
TABLE II: Comparison of multiple-camera road network extraction on nuScenes [24] dataset assessed by Landmark Precision-Recall (L-P, L-R, L-F) and Reachability Precision-Recall (R-P, R-R, R-F). ResNet-50 [57] is applied as image backbone by default. $\star$” use coupled road network. “†” use VoVNetV2 [58]
@@
AR-RNTR 62.5 47.7 54.1 73.2 52.8 61.3 0.1 (1.0$\times$)
SAR-RNTR 65.3 49.1 56.0 74.1 59.1 65.7 0.6 (6.0$\times$)
NAR-RNTR 64.1 48.3 55.1 74.0 58.7 65.1 5.5 (47$\times$)
@@
SAR-RNTR† 66.0 55.9 60.5 74.5 61.1 67.1 0.6 (6.0$\times$)
NAR-RNTR† 65.6 55.7 60.2 73.4 60.0 66.0 4.7 (47$\times$)
AR-RNTR + TIT 67.7 53.6 59.8 79.3 63.8 70.7 0.1 (1.0$\times$)
SD-RNTR + TIT 77.1 69.4 73.0 75.9 68.9 72.2 0.1 (1.0$\times$)
Figure 13: Qualitative results on nuScenes validation set. All three variants of RoadNetTransformer predict high quality road network. Only slight errors (pointed by red arrow) occur when predicting landmarks locations for AR-RNTR and NAR-RNTR.
Landmark and Reachability Table II evaluate our methods on Landmark Precision-Recall and Reachability Precision-Recall. The auto-regressive RNTR performs worst in most metrics without saying it’s slowest inference speed. In comparison, the Semi-Autoregressive RNTR leads in all metrics, and boosts the inference speed by 6.0 times. Remarkably, the non-autoregressive version of RNTR outperforms the AR-RNTR in all metrics and dramatically improves inference speed by a factor of 47$\times$, surpassing the nuScenes camera sampling frequency of 2Hz and enabling real-time inference. The small gap between NAR-RNTR and SAR-RNTR proves the feasibility of the masked sequence training and the iterative based inference. To summarize, the SAR-RNTR achieves the best performance, while the NAR-RNTR strikes a very balance between efficiency and accuracy.
TABLE III: Comparison of multiple-camera road network extraction on nuScenes [24] dataset with different split assessed by Landmark Precision-Recall (L-P, L-R, L-F) and Reachability Precision-Recall (R-P, R-R, R-F). ResNet-50 [57] is applied as image backbone by default.$\star$” use coupled road network.
@@
Over-fitting Table III shows results with the new split on the nuScenes dataset. Compared to the default split, the results from the non-overlapping split show a significant reduction in both Landmark and Reachability metrics. The decrease in Landmark performance aligns with the observation that without overfitting, invisible roads cannot be predicted through round-view images. Comparing the three variants of RoadNetTransformer, the Landmark scores show no significant differences, all hovering around 36.5, suggesting that landmark detection has reached its upper limit. In non-overlapping scenes, the primary challenge is the detection of roads, which leads to less accurate landmark detection compared to the overfitting scenario. This inaccuracy also adversely affects Reachability.
Figure 14: Qualitative results on the nuScenes [24] non-overfitting validation set. Both AR-RNTR and AR-RNTR with Topology-Inherited Training do not perform ideally in the non-overfitting case. However, with the integration of SD-Maps, AR-RNTR can finally produce high-quality and usable road network.
Topology-Inherited Training (TIT) After applying Topology-Inherited Training, AR-RNTR experiences a significant improvement in both the overfitting and non-overfitting splits. In the non-overfitting split, TIT helps AR-RNTR exceed the performance of both SAR-RNTR and NAR-RNTR. Especially in Reachability, TIT leads SAR-RNTR by 11.6, demonstrating that with the help of topology data inheritance, topology training is more effective. In terms of Landmark, however, the improvement is not as obvious, with less than a 3-point increase from AR-RNTR. This indicates that TIT cannot fundamentally improve road landmark detection.
SD-Maps The usage of SD-Maps as prior information brings significant improvements for both the overfitting and non-overfitting splits, with more pronounced benefits for the non-overfitting split. It raises the Landmark F-score from 39.1 to 54.7 and the Reachability F-score from 49.8 to 62.6. The most notable improvement in the Landmark F-score demonstrates the effectiveness of SD-Maps. Additionally, better landmark detection supports improved topology detection.
@@ IV-D Ablation studies
TABLE IV: Ablation studies on number of Transformer decoder layers and Intra-seq self-attention of NAR-RNTR. # Key-pt denotes Key-point Transformer Decoder layer number, and # Para-seq denotes Parallel-seq Transformer Decoder layer number. Intra-Seq means Intra-seq self-attention in Parallel-seq Transformer Decoder layer. VoVNetV2 [58] pretrained on extra data is applied as image backbone by default. The row with gray color is our final choice.
@@
Number of Transformer layers We investigated the impact of the number of Transformer layers on NAR-RNTR in Table IV. Using fewer Key-point Transformer decoder layers leads to a significant loss in accuracy for both landmark and reachability, but provides limited speed-up. Conversely, using fewer Parallel-seq Transformer decoder layers dramatically improves inference speed while incurring less accuracy loss.
Intra-Seq self-attention The bottom row of Table IV highlights the crucial role of Intra-seq self-attention. Its absence results in a significant loss in both landmark localization and topology connection.
TABLE V: Ablation studies on mask ratio and iteration times of NAR-RNTR. VoVNetV2 [58] pretrained on extra data is applied as image backbone by default. The row with gray color is our final choice.
@@
Mask ratio and iteration times Our investigation into the impact of the mask ratio during training of NAR-RNTR revealed the importance of using a large mask ratio. Consequently, a 90% mask ratio is used during Non-Autoregressive finetuning. We also observed accuracy saturation as the number of iterations increased. Therefore, we used 3 iterations for our approach.
TABLE VI: Ablation studies on BEV augmentation and synthetic noise objects [22] (including sequence augmentation and sequence noise padding). NAR-RNTR with VoVNetV2 [58] pretrained on extra data is applied. The row with gray color is our final choice.
@@
Non-unique Sorting We show the difference between the random ordering strategy and an ordering based on coordinates in Table
BEV augmentation The first column of Table VI shows that the BEV augmentation provides a significant 2.7/3.3 improvement on both Landmark and Reachability scores.
Synthetic noise objects The second column of Table VI shows that the sequence augmentation of Synthetic noise objects technique [22], however, leads to a drop in performance. Whereas, the third column shows that the sequence noise padding 1.6/1.7 improved on both Landmark and Reachability scores. But the sequence noise padding is less effective than BEV augmentation.
TABLE VII: Ablation studies on sequence embedding size and class weight. NAR-RNTR with VoVNetV2 [58] pretrained on extra data is applied. The row with gray color is our final choice.
@@
Class weight We exam the class weight for MLE loss, i.e., $w$ for
$\max\sum_{i=1}^{L}w_{i}\log{P(\hat{y}_{i}|y_{<i},\mathcal{F})},$ (10)
$\max\sum_{i=1}^{M}\sum_{j=1}^{L}w_{j}P(y_{i,j}\ |\ \hat{y},\mathcal{F},\mathcal{V}_{kp}),$ (11)
Due to the high frequency of Lineal for $v_{c}$ and the default index for $v_{d}$, we assign a lower weight to these categories. Although the second column of Table VII does not indicate a clear relationship between class weight and performance, using a lower weight for the loss results in more stable performance.
Embedding size If we extend the embedding size from 576 to 1000 or 2000, useless embeddings clearly harm the performance.
TABLE VIII: Ablation studies on Topology-Inherited Training. Models with ResNet [57] trained on the nuScenes [24] non-overfitting dataset are used. The ”Lidar Map” stage indicates the first stage of TIT, and ”Assembled” represents the third stage of TIT. ”Sweep data” uses 20Hz sampling rate data from the nuScenes dataset.
@@
Topology-Inherited Training In Table VIII, we present the ablation study for Topology-Inherited Training (TIT). In the first stage, we train using the lidar map as input. For the baseline, we use the same settings as AR-RNTR, but change the input to the lidar map. Since the input is simpler this time, adding more intense augmentation can significantly reduce overfitting. However, compared to AR-RNTR in Table III, the Landmark score shows a noticeable improvement due to the easier input modality. Nonetheless, the reachability performance is not as good as with the round-view image. The potential problem is overfitting with this simple modality, so we use the sweep data, which is ten times the sample dataset. After training on the sweep data, both the landmark and reachability metrics achieve the best results.
@@
IV-E Qualitative results
We present the visualization on overfitting datsaet in Figure 13. The precise localization of landmarks, accurate topological connections, and precise curve shapes demonstrate the superiority of RoadNetTransformer. Additional qualitative results are presented in the Supplementary material.
We also compare AR-RNTR, Topology-Inherited Training, and SD-Maps on the nuScenes [24] non-overfitting split. After adding Topology-Inherited Training, the landmark predictions are more precise, and the connections are reasonable in most cases. With the addition of SD-Maps, the results finally become practical for real applications.
V Conclusion
In summary, our work introduces a lossless, efficient and interactional sequence representation called RoadNet Sequence, which preserves both Euclidean and non-Euclidean data of road networks. We have designed an Auto-Regressive RoadNetTransformer as a baseline model, which takes advantage of the auto-regressive dependency of RoadNet Sequence. Additionally, we have proposed Semi-Autoregressive and Non-Autoregressive RoadNetTransformer models, which decouple the auto-regressive dependency of RoadNet Sequence, resulting in significantly faster inference speeds and improved accuracy. Our extensive experiments demonstrate the superiority of our RoadNet Sequence representation and RoadNetTransformer models.

### v3

@@ IV-A Dataset
To demonstrate the generalization ability of our approach, we further evaluate it on the Argoverse [58] dataset. The Argoverse dataset is a large-scale autonomous driving benchmark comprising diverse multimodal data collected from real-world urban driving in Pittsburgh and Miami. It contains 113 sequences with over 290,000 annotated 3D object tracks, covering approximately 5 million object instances. The dataset is complemented by high-definition maps spanning nearly 290 kilometers of road networks, which include lane geometry, drivable areas, and traffic semantics, enabling rich contextual modeling.
@@ IV-B Implementation
Pretrain. We follow LSS [13] for BEV Encoder. The input images are resized to $128\times 352$, and the target BEV area is from -48 to 48m in x-direction (roll) and -32 to 32m in y-direction (pitch) with resolution 11m in the ego coordinate system. ResNet-50 [57] or VoVNetV2 [59] are used as the image backbone. The ResNet-50 [57] model is pretrained using DeepLabV3+ [60] on the Cityscapes dataset[61], and the VoVNetV2 is pretrained on FCOS3D [62]. To be noted that the VoVNetV2 is pretrained on the default split of the nuScenes [24] dataset; therefore, we will not use the VoVNetV2 in the non-overfitting split. Initially, we pretrain the BEV encoder on centerline segmentation, following all the training strategies outlined in LSS [13]. During the training of RoadNetTransformer, we load and freeze the parameters of the BEV encoder.
RoadNetTransformer. Details of three variants are shown below. AR-RNTR: The length of RoadNet Sequence is padded to $6\times 100$. Transformer decoder layers of AR-RNTR are set as 6. Our AR-RNTR is trained on 300 epochs with learning rate $2\times 10^{-4}$, batch size as $2\times 8$. During the training process, we apply random flip, random rotation, and random scaling on BEV feature similar to [29, 16]. SAR-RNTR: Transformer decoder layers of SAR-RNTR are set as 6 (Key-point) plus 3 (Parallel-seq). We use noise vertices to pad all sub-sequences to the length of $6\times 18$, and the max number of key-points is set to 34. The strategies for data augmentation and training are the same as AR-RNTR. NAR-RNTR: During training, we load parameters of trained SAR-RNTR and then finetune it with a mask language modeling strategy for 100 extra epochs. During inference, the iteration number is set to 3.
Topology-Inherited Training. We apply Topology-Inherited Training only to AR-RNTR. The basic settings of AR-RNTR are the same as those previously introduced for RoadNetTransformer. In the first stage, we train on 20Hz data for 300 epochs. We use more data augmentation than when training RoadNetTransformer. The random rotation is increased from $45^{\circ}$ to $45^{\circ}$, and the scale is adjusted from 0.9 to 1.1. In the second stage, we train for 24 epochs. In the third stage, we train for 100 epochs.
SD-Map. We locate Singapore onenorth, Singapore Queenstown, Singapore Hollandvillage, and Boston Seaport from OpenStreetMap [27] and approximately choose the corresponding areas. After exporting the topology data, we retain only the nodes of highway and their secondary type of link roads.
Metrics. Thresholds for Mean Precision-Recall and Landmark Precision-Recall are uniformly sampled from $[0.5,5.0]$ m. Thresholds for Reachability Precision-Recall are uniformly sampled from $[0.5,2.5]$ m. We only take into account the reachability between vertices that are within a maximum of 5 edges of connection.
@@ IV-C Results
TABLE II: Comparison of front-camera road network extraction with state of the art on nuScenes [24] PON validation split [25]. ResNet-50 [57] is applied as image backbone by default. M-P, M-R, M-F stand for mean precision/recall/F1-score [17]. Detect stands for Detection ratio metrics [17]. C-P, C-R, C-F stand for connectivity precision/recall/F1-score [17].
@@
TABLE III: Comparison on Argoverse [58] dataset. ResNet-50 [57] is applied as image backbone by default. M-F and C-F stand for mean and connectivity F1-score [17]. Detect stands for detection ratio metrics [17].
Methods M-F Detect C-F
PINET [63] 47.2 15.1 -
TR [17] 55.6 60.1 54.9
TPLR [18] 57.1 64.2 58.1
Ours (AR-RNTR) 57.6 64.5 58.2
Comparison with state of the art. We compare our model with previous state-of-the-art methods on high-definition road network topology extraction. To achieve a fair comparison, we only utilize front-view images as input and ResNet-50 [57] trained on ImageNet-1K [64] as the backbone. Also, we utilize the PON nuScenes train/val split [25] as also applied in [17, 18]. Table II presents the results of our approach on the nuScenes dataset using the Mean Precision-Recall, Detection, and Connectivity metrics [17]. On the one hand, our model outperforms all three variants across all metrics, demonstrating the superior performance of RNTR in both centerline detection and centerline association estimation. This remarkable improvement on both centerline location and connectivity can be attributed to the unified representation provided by RoadNet Sequence and the global context reasoning capabilities of the Transformer architecture. We also validate our method on Argoverse dataset in Table III. The experimental results demonstrate that our method consistently outperforms the competitors.
On the other hand, unlike in Natural Language Processing, our Semi-Autoregressive and Non-Autoregressive versions significantly outperform the Auto-Regressive version, highlighting our dependency decoupling. sc
TABLE IV: Comparison of multiple-camera road network extraction on nuScenes [24] dataset assessed by Landmark Precision-Recall (L-P, L-R, L-F) and Reachability Precision-Recall (R-P, R-R, R-F). ResNet-50 [57] is applied as image backbone by default. $\star$” use coupled road network. “†” use VoVNetV2 [59]
@@
[…]
@@
Over-fitting. Table V shows results with the new split on the nuScenes dataset. Compared to the default split, the results from the non-overlapping split show a significant reduction in both Landmark and Reachability metrics. The decrease in Landmark performance aligns with the observation that without overfitting, invisible roads cannot be predicted through round-view images. Comparing the three variants of RoadNetTransformer, the Landmark scores show no significant differences, all hovering around 36.5, suggesting that landmark detection has reached its upper limit. In non-overlapping scenes, the primary challenge is the detection of roads, which leads to less accurate landmark detection compared to the overfitting scenario. This inaccuracy also adversely affects Reachability.
Topology-Inherited Training (TIT.) After applying Topology-Inherited Training, AR-RNTR experiences a significant improvement in both the overfitting and non-overfitting splits. In the non-overfitting split, TIT helps AR-RNTR exceed the performance of both SAR-RNTR and NAR-RNTR. Especially in Reachability, TIT leads SAR-RNTR by 11.6, demonstrating that with the help of topology data inheritance, topology training is more effective. In terms of Landmark, however, the improvement is not as obvious, with less than a 3-point increase from AR-RNTR. This indicates that TIT cannot fundamentally improve road landmark detection.
SD-Maps. The usage of SD-Maps as prior information brings significant improvements for both the overfitting and non-overfitting splits, with more pronounced benefits for the non-overfitting split. It raises the Landmark F-score from 39.1 to 54.7 and the Reachability F-score from 49.8 to 62.6. The most notable improvement in the Landmark F-score demonstrates the effectiveness of SD-Maps. Additionally, better landmark detection supports improved topology detection.
@@ IV-D Ablation Studies
TABLE VI: Ablation studies on number of Transformer decoder layers and Intra-seq self-attention of NAR-RNTR. # Key-pt denotes Key-point Transformer Decoder layer number, and # Para-seq denotes Parallel-seq Transformer Decoder layer number. Intra-Seq means Intra-seq self-attention in Parallel-seq Transformer Decoder layer. VoVNetV2 [59] pretrained on extra data is applied as image backbone by default. The row with gray color is our final choice.
@@
Number of Transformer layers. We investigated the impact of the number of Transformer layers on NAR-RNTR in Table VI. Using fewer Key-point Transformer decoder layers leads to a significant loss in accuracy for both landmark and reachability, but provides a limited speedup. Conversely, using fewer Parallel-seq Transformer decoder layers dramatically improves inference speed while incurring less accuracy loss.
Intra-Seq self-attention. The bottom row of Table VI highlights the crucial role of Intra-seq self-attention. Its absence results in a significant loss in both landmark localization and topology connection.
TABLE VII: Ablation studies on mask ratio and iteration times of NAR-RNTR. VoVNetV2 [59] pretrained on extra data is applied as image backbone by default. The row with gray color is our final choice.
@@
Mask ratio and iteration times. Our investigation into the impact of the mask ratio during the training of NAR-RNTR revealed the importance of using a large mask ratio. Consequently, a 90% mask ratio is used during Non-Autoregressive finetuning. We also observed accuracy saturation as the number of iterations increased. Therefore, we used 3 iterations for our approach.
TABLE VIII: Ablation studies on BEV augmentation and synthetic noise objects [22] (including sequence augmentation and sequence noise padding). NAR-RNTR with VoVNetV2 [59] pretrained on extra data is applied. The row with gray color is our final choice.
@@
TABLE IX: Ablation studies on sequence embedding size and class weight. NAR-RNTR with VoVNetV2 [59] pretrained on extra data is applied. The row with gray color is our final choice.
@@
TABLE X: Ablation studies on Topology-Inherited Training. Models with ResNet [57] trained on the nuScenes [24] non-overfitting dataset are used. The ”Lidar Map” stage indicates the first stage of TIT, and ”Assembled” represents the third stage of TIT. ”Sweep data” uses 20Hz sampling rate data from the nuScenes dataset.
@@
TABLE XI: Comparison on nuScenes [24] dataset assessed by Landmark Precision-Recall (L-P, L-R, L-F) and Reachability Precision-Recall (R-P, R-R, R-F). The model used is AR-RNTR. ResNet-50 [57] is applied as image backbone by default.
Learning Rate Landmark Reachability
L-P L-R L-F R-P R-R R-F
1e-4 61.0 46.8 53.0 72.5 52.1 60.6
2e-4 62.5 47.7 54.1 73.2 52.8 61.3
3e-4 62.0 47.2 54.0 72.8 52.4 60.7
TABLE XII: Comparison on nuScenes [24] dataset assessed by Landmark Precision-Recall (L-P, L-R, L-F) and Reachability Precision-Recall (R-P, R-R, R-F). ResNet-50 [57] is applied as image backbone by default.
Methods Landmark Reachability
L-P L-R L-F R-P R-R R-F
[…]
TABLE XIII: Comparison of different geographical location settings on nuScenes [24] PON validation split [25]. ResNet-50 [57] is applied as image backbone by default. M-P, M-R, M-F stand for mean precision/recall/F1-score [17]. Detect stands for Detection ratio metrics [17]. C-P, C-R, C-F stand for connectivity precision/recall/F1-score [17].
Method nuScenes
Singapore Boston
M-P M-R M-F Detect C-P C-R C-F M-P M-R M-F Detect C-P C-R C-F
AR-RNTR 61.1 57.8 59.6 61.8 63.1 52.5 57.9 60.8 58.2 59.1 61.5 63.4 52.8 57.3
SAR-RNTR 63.8 60.4 62.0 63.4 67.3 56.8 61.2 63.0 59.6 61.1 63.5 66.8 57.5 62.0
NAR-RNTR 61.4 59.1 60.2 62.3 66.8 55.7 60.5 62.5 59.6 61.1 61.8 66.2 56.4 61.2
TABLE XIV: Comparison of different geographical location settings on Argoverse [58] dataset. ResNet-50 [57] is applied as image backbone by default. M-F and C-F stand for mean and connectivity F1-score [17]. Detect stands for detection ratio metrics [17].
Method Argoverse
Miami Pittsburgh
M-F Detect C-F M-F Detect C-F
AR-RNTR 58.2 64.8 58.4 57.2 64.1 58.1
SAR-RNTR 60.1 65.4 61.2 60.6 66.3 62.0
NAR-RNTR 58.4 65.6 60.7 57.6 64.3 60.2
TABLE XV: Comparison of different weather condition settings on nuScenes [24] PON validation split [25]. ResNet-50 [57] is applied as image backbone by default. M-F stand for mean F1-score [17]. Detect stands for Detection ratio metrics [17]. C-F stand for connectivity F1-score [17].
Method nuScenes
Normal Rainy Cloudy
M-F Detect C-F M-F Detect C-F M-F Detect C-F
AR-RNTR 59.9 62.8 58.4 58.6 61.3 56.9 59.0 61.5 57.3
SAR-RNTR 62.4 63.9 62.6 60.7 63.1 61.1 61.8 63.6 62.0
NAR-RNTR 62.0 63.5 62.1 60.3 61.5 60.3 60.2 61.7 60.4
Non-unique sorting. We show the difference between the random ordering strategy and an ordering based on coordinates in Table
BEV augmentation. The first column of Table VIII shows that the BEV augmentation provides a significant 2.7/3.3 improvement in both Landmark and Reachability scores.
Synthetic noise objects. The second column of Table VIII shows that the sequence augmentation of Synthetic noise objects technique [22], however, leads to a drop in performance. Whereas, the third column shows that the sequence noise padding 1.6/1.7 improved on both Landmark and Reachability scores. But the sequence noise padding is less effective than BEV augmentation.
Class weight. We exam the class weight for MLE loss, i.e., $w$ for
$\max\sum_{i=1}^{L}w_{i}\log{P(\hat{y}_{i}|y_{<i},\mathcal{F})}.$ (13)
$\max\sum_{i=1}^{M}\sum_{j=1}^{L}w_{j}P(y_{i,j}\ |\ \hat{y},\mathcal{F},\mathcal{V}_{kp}).$ (14)
Due to the high frequency of Lineal for $v_{c}$ and the default index for $v_{d}$, we assign a lower weight to these categories. Although the second column of Table IX does not indicate a clear relationship between class weight and performance, using a lower weight for the loss results in more stable performance.
Embedding size. If we extend the embedding size from 576 to 1000 or 2000, useless embeddings clearly harm the performance.
Topology-Inherited Training. In Table X, we present the ablation study for Topology-Inherited Training (TIT). In the first stage, we train using the lidar map as input. For the baseline, we use the same settings as AR-RNTR, but change the input to the lidar map. Since the input is simpler this time, adding more intense augmentation can significantly reduce overfitting. However, compared to AR-RNTR in Table V, the Landmark score shows a noticeable improvement due to the easier input modality. Nonetheless, the reachability performance is not as good as with the round-view image. The potential problem is overfitting with this simple modality, so we use the sweep data, which is ten times the sample dataset. After training on the sweep data, both the landmark and reachability metrics achieve the best results.
@@
Parameter fine-tuning. We conducted experiments on hyperparameter tuning by varying the learning rate. The results, presented in Table XI, show that our model performs best with the default setting and remains stable with minimal variation when the learning rate is adjusted.
Computational costs. Besides the FPS reported in Table IV, we further provide details on the model parameters and inference time. The base model, AR-RNTR, has 37.25 million parameters, while NAR-RNTR has 35.90 million parameters and requires approximately 6GB of memory during inference with a batch size of 2, where the input images are 128 × 352.
Standard deviations. Including standard deviations strengthens the reliability of our findings. We have provided the mean and standard deviation for AR-RNTR, SAR-RNTR, and NAR-RNTR, along with the number of trials. Each experiment was repeated five times, and the results are presented in Table XII. The performance of our models is statistically stable, with minimal variation.
Figure 16: Mean Precision/Recall v.s. thresholds. $\dagger$}” use VoVNetV2 [59] pretrained on extra data as backbone. Thresholds are from [0.5,1.0,1.5,2.0,2.5,3.0,3.5,4.0,4.5,5.0]m[0.5,1.0,1.5,2.0,2.5,3.0,3.5,4.0,4.5,5.0]m.
IV-E Precision-Recall Curve
IV-F Robustness
To evaluate the robustness of our method, we group the evaluation results by geographic region and report performance separately. As shown in Table XIII and Table XIV, our model consistently performs well across diverse urban environments. Specifically, in the nuScenes dataset, we observe strong and stable performance in both Singapore and Boston, demonstrating generalization across cities with different road structures and traffic styles. Similarly, on the Argoverse dataset, which contains data from Miami and Pittsburgh, our method also shows solid performance, indicating good adaptability to diverse North American cities.
Furthermore, to assess robustness under varying weather conditions, we conduct additional experiments on the nuScenes validation split by categorizing scenes based on annotated weather labels (i.e., normal, rainy, and cloudy). As reported in Table XV, our method maintains strong performance across all weather conditions. This confirms the robustness and reliability of our model under challenging real-world scenarios.
Overall, the results across these three tables clearly demonstrate that our approach is resilient to variations in both geography and weather, supporting its effectiveness in real-world deployment.
V Conclusion and Discussions
Conclusion. In summary, our work introduces a lossless, efficient, and interactional sequence representation called RoadNet Sequence, which preserves both Euclidean and non-Euclidean data of road networks. We have designed an Auto-Regressive RoadNetTransformer as a baseline model, which takes advantage of the auto-regressive dependency of RoadNet Sequence. Additionally, we have proposed Semi-Autoregressive and Non-Autoregressive RoadNetTransformer models, which decouple the auto-regressive dependency of RoadNet Sequence, resulting in significantly faster inference speeds and improved accuracy. Our extensive experiments demonstrate the superiority of our RoadNet Sequence representation and RoadNetTransformer models.